OpenStudy (anonymous):

If sin a=.25 and cos b=.25 what is sin b/2 + cos a/2?

OpenStudy (anonymous):

is 1.604680374

OpenStudy (anonymous):

I'm getting 2 ???

OpenStudy (anonymous):

probably not????

OpenStudy (anonymous):

i'm sorry

OpenStudy (anonymous):

0.375

OpenStudy (anonymous):

1.343245

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

No problem literally this is all GREEK to me!! :D appreciate your help

OpenStudy (anonymous):

\[\sin ^{2}\left( \frac{ b }{ 2 } \right) = \frac{ 1 - \cos b }{ 2 }\]and\[\cos ^{2}\left( \frac{ a }{ 2 } \right) = \frac{ 1 + \cos a }{ 2 }\]where\[\cos a = \sqrt{1 - \sin ^{2}a}\]Using that last equation,\[\cos a = \sqrt{1 - (0.25)^{2}} = 0.9682458\]Using this in the second equation\[\cos ^{2}\left( \frac{ a }{ 2 } \right) = \frac{ 1 + 0.9682458 }{ 2 } = 0.9841229\]

OpenStudy (anonymous):

The first equation is easy.\[\sin ^{2}\left( \frac{ b }{ 2 } \right) = \frac{ 1 - 0.25 }{ 2 } = 0.375\]So, now you just add the sqrt of this and the sqrt from my previous post.

OpenStudy (anonymous):

1.604329687

OpenStudy (anonymous):

Yes, that's close enough. I rounded when I gave you some of the partial results from calculations. btw, was this a multiple-choice question?

OpenStudy (anonymous):

yes it was thank you so so so much!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Well, between Openstudy MathJax not working for me for a while, and my mistake from before, I'm surprised you waited for me!

OpenStudy (anonymous):

What did your multiple-choice answers look like exactly? That would be very informative.

OpenStudy (anonymous):

Needed your help badly wrote a comment for you " GOOD ONE!"

OpenStudy (anonymous):

1.60 ,1.32,.26,1.06

OpenStudy (anonymous):

Ok, that's good. What I was looking for was whether or not the answer contained radical signs and if it didn't, how far it went with the decimal places. It confirms the validity of my method, although I'm going to look at this one for a while and see if there was an easier way. If you don't stick around, I'll message you on what I find.

OpenStudy (anonymous):

Thank you wasn't given any formulas to solve from class been looking on internet for trig formula sheets

OpenStudy (anonymous):

So Bright makes Sheldon from " Big Bang Theory" look ignorant!!!!!! posted as a testimony for you !! Thanks man

OpenStudy (anonymous):

OK! I just went through it all again, and if there is any shorter way, it cannot be much shorter at all. I really think you have to go through all these steps to get the answer, and it's obviously a lot of work. But it's fun! And I just saw your Sheldon comment. It's not true of course, but maybe I should start wearing long sleeves under my t-shirts! :-) Anyway, this was a lot of fun (for math nerds like me!) And you're wonderful to work with! Thanks for the very nice comments!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are very very welcome! @maltman0004 I had to look at your profile to realize that you are Melissa. Before I did that, I thought you were a guy that likes malts!

OpenStudy (anonymous):

lol its my college ID MALTMAN lol last name Altman and 1st initial

OpenStudy (anonymous):

In any case, I like malts, I like Sheldon, and I sure like working with you! @maltman0004